# Human-in-the-Loop Generation of Adversarial Texts: A Case Study on Tibetan Script

### v3

@@ Abstract
DNN-based language models perform excellently on various tasks, but even SOTA LLMs are susceptible to textual adversarial attacks. Adversarial texts play crucial roles in multiple subfields of NLP. However, current research has the following issues. (1) Most textual adversarial attack methods target rich-resourced languages. How do we generate adversarial texts for less-studied languages? (2) Most textual adversarial attack methods are prone to generating invalid or ambiguous adversarial texts. How do we construct high-quality adversarial robustness benchmarks? (3) New language models may be immune to part of previously generated adversarial texts. How do we update adversarial robustness benchmarks? To address the above issues, we introduce HITL-GAT [^1] Video Demonstration:
https://youtu.be/xFrto00rHuI
@@
https://github.com/metaphors/HITL-GAT
@@
, a system based on a general approach to human-in-the-loop generation of adversarial texts. Additionally, we utilize HITL-GAT to make a case study on Tibetan script which can be a reference for the adversarial research of other less-studied languages.
@@ A Case Study on Tibetan Script
Xi Cao♡♠,  Yuan Sun♡♠🖂, Jiajun Li♢♣,  Quzong Gesang♢♣,  Nuo Qun♢♣🖂,  Tashi Nyima♢♣ National Language Resource Monitoring & Research Center Minority Languages Branch, Beijing, China Minzu University of China, Beijing, China Collaborative Innovation Center for Tibet Informatization, Lhasa, China Tibet University, Lhasa, China caoxi@muc.edu.cn, sunyuan@muc.edu.cn, q_nuo@utibet.edu.cn
@@ 1 Introduction
Figure 1: Workflow of HITL-GAT. While a new foundation model, downstream dataset, or textual adversarial attack method emerges, we can enter the loop to make the adversarial robustness benchmark evolve.
The vulnerability of DNNs to adversarial attacks was first identified in CV (Szegedy et al., 2014; Goodfellow et al., 2015). The adversarial attack refers to an attack method in which the attacker adds imperceptible perturbations to the original input, resulting in the incorrect judgment of a DNN. Later, NLP researchers found NLP applications based on DNNs are also vulnerable to adversarial attacks (Jia and Liang, 2017; Ebrahimi et al., 2018a, b). The examples generated during textual adversarial attacks are called adversarial texts. Adversarial texts play crucial roles in multiple subfields of NLP (Chen et al., 2022). In the security field, adversarial texts can reveal the robustness shortcomings of NLP models; In the explainability field, adversarial texts can partly explain the decision process of NLP models; In the evaluation field, adversarial robustness benchmarks can stress-test the comprehension of NLP models; In the data augmentation field, adversarial training can improve the performance and robustness of NLP models.
@@
Currently, textual adversarial attack methods with different granularity (e.g. character-, word-, and sentence-level), in different settings (white- and black-box), and for different tasks (text classification, text generation, etc.) have been proposed (Goyal et al., 2023). Due to the general adaptability of models to the classification task, adversarial robustness evaluation is mainly focused on this task. Additionally, most of the methods target rich-resourced languages, especially English. However, because of the differences in language resources and textual features, it is challenging to transfer these methods to other languages. Problem 1: How do we generate adversarial texts for less-studied languages?
@@
Wang et al. (2021a) apply 14 textual adversarial attack methods to GLUE tasks (Wang et al., 2019) to construct the widely used adversarial robustness benchmark AdvGLUE. In their construction, they find that most textual adversarial attack methods are prone to generating invalid or ambiguous adversarial texts, with around 90% either changing the original semantics or hindering the annotators’ unanimity. In our case study on Tibetan script, we also come to the same conclusion. Problem 2: How do we construct high-quality adversarial robustness benchmarks?
@@
Wang et al. (2023) employ ANLI (Nie et al., 2020) and AdvGLUE (Wang et al., 2021a) to assess the adversarial robustness of ChatGPT and several previous popular foundation models and find ChatGPT is the best. However, both ANLI and AdvGLUE are constructed using fine-tuned BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019) as victim models. Language models are evolving, while adversarial robustness benchmarks never. We argue that new language models may be immune to part of previously generated adversarial texts. Less-studied languages are at a very early stage of adversarial robustness evaluation compared to rich-resourced languages, and it is essential to envisage sustainable adversarial robustness evaluation in advance. Problem 3: How do we update adversarial robustness benchmarks?
@@
To address the above problems, we introduce HITL-GAT, a system for human-in-the-loop generation of adversarial texts. Figure 1 depicts the workflow of HITL-GAT. In a loop where a new foundation model, downstream dataset, or textual adversarial attack method emerges, our team starts to construct victim models, generate adversarial examples, construct high-quality benchmarks, and evaluate adversarial robustness. The loop allows adversarial robustness benchmarks to evolve along with new models, datasets, and attacks (Problem 3). Figure 2 depicts the four stages in one pipeline detailedly. Firstly, we fine-tune the previous model and the new model on the same downstream datasets to construct victim models. Subsequently, we implement adversarial attacks on the victim models constructed from the previous model upon downstream datasets to generate adversarial examples. Afterward, we customize filter conditions and conduct human annotation to construct a high-quality adversarial robustness benchmark (Problem 2). Finally, we evaluate the adversarial robustness of the new model on the benchmark. Additionally, we make a case study on one less-studied language, Tibetan script, based on the general human-in-the-loop approach to adversarial text generation (Problem 1).
@@
(2) We develop an interactive system called HITL-GAT based on the approach to human-in-the-loop generation of adversarial texts. This system is successfully applied to a case study on one less-studied language.
(3) We utilize HITL-GAT to make a case study on Tibetan script and construct the first adversarial robustness benchmark for Tibetan script called AdvTS under the existing conditions. This case study can be a reference for the adversarial research of other less-studied languages.
(4) We open-source both the system and the case study to facilitate future explorations.
@@ 2.1 Textual Adversarial Attack Frameworks
TextAttack (Morris et al., 2020) and OpenAttack (Zeng et al., 2021) are two powerful and easy-to-use Python frameworks for textual adversarial attacks. They are both for text classification, with similar toolkit functionality and complementary attack methods. From a developer’s perspective, TextAttack utilizes a relatively rigorous architecture to unify different attack methods, while OpenAttack is more flexible. SeqAttack (Simoncini and Spanakis, 2021) and RobustQA (Boreshban et al., 2023) are textual adversarial attack frameworks for named entity recognition and question answering, respectively. These frameworks provide an excellent platform for customizing textual adversarial attack methods to stress-test the adversarial robustness of NLP models automatically.
@@ 2.2 Human-in-the-Loop Adversarial Text Generation
Most goals of using a human-in-the-loop approach in NLP tasks are to improve the model performance in various aspects (Wang et al., 2021b). With these goals, language models evolve. As continuous advancement of model capabilities, it is imperative to explore the paradigm for benchmark evolution.
Wallace et al. (2019) guide human authors to keep crafting adversarial questions to break the question answering models with the aid of visual model predictions and interpretations. They conduct two rounds of adversarial writing. In the first round, human authors attack a traditional ElasticSearch model A to construct the adversarial set x. Then, they use x to evaluate A, a bidirectional recurrent neural network model B, and a deep averaging network model C. In the second round, they train A, B, and C on a larger dataset. Human authors attack A and B to construct the adversarial set x and x’. Then, they use x and x’ to evaluate A, B, and C. We see their human-in-the-loop approach as an embryo of adversarial robustness benchmark evolution, although with high labor costs.
Wang et al. (2021a) leverage the automation of textual adversarial attack methods as well as metric and human filtering to construct the adversarial robustness benchmark AdvGLUE, which is widely used from the BERT period (Wang et al., 2021a) to the ChatGPT period (Wang et al., 2023). On the one hand, the results show that the model is progressively having stronger robustness; But on the other hand, it also suggests that the benchmark is gradually becoming outdated.
@@ 3 Implementation
3.1 Definition
Due to the general adaptability of language models to the text classification task, our work focuses on the adversarial robustness evaluation of language models on this task. The definition of textual adversarial attacks on text classification is as follows.
For a text classifier $F$, let $x$ ($x\in{X}$, $X$ includes all possible input texts) be the original input text and $y$ ($y\in{Y}$, $Y$ includes all possible output labels) be the corresponding output label of $x$, denoted as
$F(x)={\mathop{\arg\max}_{\dot{y}\in{Y}}{P(\dot{y}|x)}}={y}.$
For a successful textual adversarial attack, let $x^{\prime}=x+\delta$ be the perturbed input text, where $\delta$ is the imperceptible perturbation, denoted as
$F(x^{\prime})={\mathop{\arg\max}_{\dot{y}\in{Y}}{P(\dot{y}|x^{\prime})}}\neq{y}.$
Figure 3: Screenshots of HITL-GAT.
3.2 Overview
Our system for human-in-the-loop generation of adversarial texts, HITL-GAT, contains four stages in one pipeline: victim model construction, adversarial example generation, high-quality benchmark construction, and adversarial robustness evaluation. Figure 2 depicts the flowchart of HITL-GAT. These four stages will be detailed in the following four subsections respectively. Our flexible interactive system allows users to either go through the entire pipeline or directly start at any stage.
Gradio (Abid et al., 2019) is an open-sourced Python package that allows developers to quickly build a web demo or application for machine learning. LlamaBoard is the user-friendly GUI (Graphical User Interface) of LlamaFactory (Zheng et al., 2024). The GUI of our system is powered by Gradio and draws inspiration from the design of LlamaBoard. Figure 3 shows the screenshots of HITL-GAT.
@@ 3.3 Construct Victim Models
When a new foundation model B emerges, in order to better evaluate the adversarial robustness of B, we need to quantitatively and thoroughly perform evaluation on multiple downstream tasks. For the purpose of stress-testing the adversarial robustness of B more effectively, i.e., constructing a stronger adversarial robustness benchmark with high quality, we can choose at least one previous SOTA or similar-structured foundation model A to implement textual adversarial attacks on it to generate updated adversarial texts. We can also follow this stage when a new downstream dataset n is available.
@@ 3.4 Generate Adversarial Examples
The way human authors keep writing adversarial texts (Wallace et al., 2019) is high-labor-cost. With the emergence of different textual adversarial attack methods, such as TextBugger (Li et al., 2019), TextFooler (Jin et al., 2020), BERT-ATTACK (Li et al., 2020), SememePSO-Attack (Zang et al., 2020), and SemAttack (Wang et al., 2022), adversarial texts generation has become relatively easy. Due to the out-of-the-box and extensible features of textual adversarial attack frameworks, such as TextAttack (Morris et al., 2020) and OpenAttack (Zeng et al., 2021), for rich-resourced languages, especially English, the acquisition of attack methods is low-cost; for less-studied languages, the customization of attack methods is additionally necessary. We can directly enter this stage when a new textual adversarial attack N appears.
In this stage, we implement textual adversarial attacks I,II,...,N on the victim language models constructed from foundation model A upon the test set of downstream datasets 1,2,...,n to generate the first-round adversarial texts automatically. The adversarial example generation stage is depicted in the second part of Figure 2.
@@ 3.5 Construct High-Quality Benchmarks
The construction process of AdvGLUE (Wang et al., 2021a), a widely used adversarial robustness benchmark, tells us that most textual adversarial attack methods are prone to generating invalid or ambiguous adversarial texts, with around 90% either changing the original semantics or hindering the annotators’ unanimity. Therefore, human annotation is indispensable and can make benchmarks more practical and relevant. In order to reduce the cost of human annotation, the first-round adversarial texts need to be screened automatically first using appropriate filter conditions. Due to the fact that humans perceive texts through their eyes and brains, both filter conditions and human annotation should follow the visual and semantic similarity between adversarial texts and original texts. Filter conditions can be the following metrics: Edit Distance, Normalized Cross-Correlation Coefficient (from the perspective of visual similarity); Cosine Similarity, BERTScore (Zhang et al., 2020) (from the perspective of semantic similarity); and so on. Human annotation still requires additional consideration of annotators’ unanimity so that adversarial texts can be deemed human-acceptable. For example, given an original text and an adversarial text, we ask several annotators to score the human acceptance of the adversarial text based on the visual and semantic similarity between the two texts, from 1 to 5. The higher the score, the higher the human acceptance. If all annotators score the human acceptance of the adversarial text as 4 or 5, the adversarial text will be included in the adversarial robustness benchmark.
@@ 3.6 Evaluate Adversarial Robustness
This stage aims at quantitatively and thoroughly evaluating the adversarial robustness of new foundation models using the constructed high-quality adversarial robustness benchmark.
The adversarial robustness benchmark x is a collection of $n$ subsets, each of which contains high-quality adversarial texts generated from the test set of the corresponding downstream dataset. We take the average accuracy on $n$ subsets as the adversarial robustness ($AdvRobust$) of the new foundation model B on x, denoted as:
@@
In this stage, we utilize the constructed high-quality adversarial robustness benchmark x to evaluate the adversarial robustness of the foundation model B quantitatively and thoroughly. The adversarial robustness evaluation stage is depicted in the fourth part of Figure 2.

### v5

@@ Abstract
DNN-based language models excel across various NLP tasks but remain highly vulnerable to textual adversarial attacks. While adversarial text generation is crucial for NLP security, explainability, evaluation, and data augmentation, related work remains overwhelmingly English-centric, leaving the problem of constructing high-quality and sustainable adversarial robustness benchmarks for lower-resourced languages both difficult and understudied. First, method customization for lower-resourced languages is complicated due to linguistic differences and limited resources. Second, automated attacks are prone to generating invalid or ambiguous adversarial texts. Last but not least, language models continuously evolve and may be immune to parts of previously generated adversarial texts. To address these challenges, we introduce HITL-GAT [^1] Video Demonstration:
https://youtu.be/tXla4yAggwA
@@
https://github.com/CMLI-NLP/HITL-GAT
@@
, an interactive system based on a general approach to human-in-the-loop generation of adversarial texts. Additionally, we demonstrate the utility of HITL-GAT through a case study on Tibetan script, employing three customized adversarial text generation methods and establishing its first adversarial robustness benchmark, providing a valuable reference for other lower-resourced languages.
@@ A Case Study on Tibetan Script
Xi Cao♡♠,  Yuan Sun♡♠🖂, Jiajun Li♢♣,  Quzong Gesang♢♣,  Nuo Qun♢♣🖂,  Tashi Nyima♢♣ Institute of National Security, Minzu University of China, Beijing, China School of Information Science and Technology, Xizang University, Lhasa, China National Language Resource Monitoring & Research Center | Minority Languages Branch, Beijing, China The State Key Laboratory of Tibetan Intelligence, Lhasa, China caoxi@muc.edu.cn, sunyuan@muc.edu.cn, q_nuo@utibet.edu.cn
@@ 1 Introduction
The adversarial attack refers to an attack method in which the attacker adds imperceptible perturbations to the original input, resulting in the incorrect judgment of a DNN-based model. The examples generated during textual adversarial attacks are called adversarial texts.
@@
Figure 1: Workflow of HITL-GAT. While a new language model, downstream dataset, or textual adversarial attack method emerges, we can enter the loop to make the adversarial robustness benchmark evolve.
@@
Due to the general adaptability of language models to classification tasks, adversarial robustness evaluation is mainly focused on the domain. Currently, most of the adversarial text generation methods target higher-resourced languages, especially English. Because of the differences in textual features and language resources, it is challenging to transfer these methods to other languages. Problem 1: How do we generate adversarial texts for lower-resourced languages?
@@
NEURIPS_DATASETS_AND_BENCHMARKS2021_335f5352 apply 14 textual adversarial attack methods to GLUE tasks (DBLP:conf/iclr/WangSMHLB19) to construct the widely used adversarial robustness benchmark AdvGLUE. In their construction, they find that most textual adversarial attack methods are prone to generating invalid or ambiguous adversarial texts, with around 90% either changing the original semantics or hindering the annotators’ unanimity. In our case study on Tibetan script, we also come to the same conclusion. Problem 2: How do we construct high-quality adversarial robustness benchmarks?
@@
wang2023on employ ANLI (nie-etal-2020-adversarial) and AdvGLUE (NEURIPS_DATASETS_AND_BENCHMARKS2021_335f5352) to assess the adversarial robustness of ChatGPT and several previous popular language models and find ChatGPT is the best. However, both ANLI and AdvGLUE are constructed using fine-tuned BERT (devlin-etal-2019-bert) and RoBERTa (liu2019robertarobustlyoptimizedbert) as victim models. Language models are evolving, while adversarial robustness benchmarks never. We argue that new language models may be immune to part of previously generated adversarial texts. Lower-resourced languages are at a very early stage of adversarial robustness evaluation compared to higher-resourced languages, and it is essential to envisage sustainable adversarial robustness evaluation in advance. Problem 3: How do we update adversarial robustness benchmarks?
@@
To address the above problems, we introduce HITL-GAT, an interactive system for human-in-the-loop generation of adversarial texts. Figure 1 depicts the workflow of HITL-GAT. In a loop where a new language model, downstream dataset, or textual adversarial attack method emerges, our team starts to construct victim models, generate adversarial examples, construct high-quality benchmarks, and evaluate adversarial robustness. The loop allows adversarial robustness benchmarks to evolve along with new models, datasets, and attacks (Problem 3). Figure 2 depicts the four stages in one pipeline detailedly. Firstly, we fine-tune the previous model and the new model on the same downstream datasets to construct victim models. Subsequently, we implement adversarial attacks on the victim models constructed from the previous model upon downstream datasets to generate adversarial examples. Afterward, we customize filter conditions and conduct human annotation to construct a high-quality adversarial robustness benchmark (Problem 2). Finally, we evaluate the adversarial robustness of the new model on the benchmark. Additionally, we make a case study on one lower-resourced language, Tibetan, based on the general human-in-the-loop approach to adversarial text generation (Problem 1).
@@
(2) We develop an interactive system called HITL-GAT based on the general approach to human-in-the-loop generation of adversarial texts. This system is successfully applied to a case study on one lower-resourced language.
(3) We demonstrate the utility of HITL-GAT through a case study on Tibetan script, employing three customized adversarial text generation methods and establishing its first adversarial robustness benchmark, providing a valuable reference for other lower-resourced languages.
(4) We open-source both the system and the case study under GNU General Public License v3.0 to facilitate future explorations. Our code repository received 42 stars, and our 12 victim models were downloaded more than 5,000 times before paper submission on November 15, 2025.
@@ 2.1 Textual Adversarial Attack Frameworks
TextAttack (morris-etal-2020-textattack) and OpenAttack (zeng-etal-2021-openattack) are two powerful and easy-to-use Python frameworks for textual adversarial attacks. They are both for text classification, supporting English and Chinese, with similar toolkit functionality and complementary attack methods. From a developer’s perspective, TextAttack utilizes a relatively rigorous architecture to unify different attack methods, while OpenAttack is more flexible. SeqAttack (simoncini-spanakis-2021-seqattack) and RobustQA (boreshban-etal-2023-robustqa) are textual adversarial attack frameworks for named entity recognition and question answering, respectively, supporting English only. These frameworks provide an excellent platform to stress-test the adversarial robustness of models targeting higher-resourced languages. However, the weaponization of lower-resourced languages against NLP security (lent2025weaponizationnlpsecuritymedium; yoo-etal-2025-code; lent-etal-2025-nlp) highlights the urgent need for research in this area. To our knowledge, HITL-GAT is the first interactive system to build adversarial robustness benchmarks from scratch for a truly low-resource language.
@@ 2.2 Human-in-the-Loop Adversarial Text Generation
wallace-etal-2019-trick guide human authors to keep crafting adversarial questions to break the question answering models with the aid of visual model predictions and interpretations. They conduct two rounds of adversarial writing. In the first round, human authors attack a traditional ElasticSearch model A to construct the adversarial set x. Then, they use x to evaluate A, a bidirectional recurrent neural network model B, and a deep averaging network model C. In the second round, they train A, B, and C on a larger dataset. Human authors attack A and B to construct the adversarial set x and x’. Then, they use x and x’ to evaluate A, B, and C. We see their human-in-the-loop approach as an embryo of adversarial robustness benchmark evolution, despite the high labor cost of relying on human authors to think and write adversarial texts. Most goals of using a human-in-the-loop approach in NLP tasks are to improve the model performance in various aspects (wang-etal-2021-putting). With these goals, language models evolve. As continuous advancement of model capabilities, it is imperative to explore the paradigm for benchmark evolution. To our knowledge, even though our work is preliminary, we are the first to explore the evolution of adversarial robustness benchmarks.
@@ 3 Implementation
Definition
Due to the general adaptability of language models to the text classification task, our work focuses on the adversarial robustness evaluation of language models on this task. The definition of textual adversarial attacks on text classification is as follows. For a text classifier $F$, let $x$ ($x\in{X}$, $X$ includes all possible input texts) be the original input text and $y$ ($y\in{Y}$, $Y$ includes all possible output labels) be the corresponding output label of $x$, denoted as $F(x)={\mathop{\arg\max}_{\dot{y}\in{Y}}{P(\dot{y}|x)}}={y}$. For a successful textual adversarial attack, let $x^{\prime}=x+\delta$ be the perturbed input text, where $\delta$ is the imperceptible perturbation, denoted as $F(x^{\prime})={\mathop{\arg\max}_{\dot{y}\in{Y}}{P(\dot{y}|x^{\prime})}}\neq{y}$.
Overview
Our system for human-in-the-loop generation of adversarial texts, HITL-GAT, contains four stages in one pipeline: victim model construction, adversarial example generation, high-quality benchmark construction, and adversarial robustness evaluation. Figure 2 depicts the flowchart of HITL-GAT. These four stages will be detailed in the following four subsections respectively. Our flexible interactive system allows users to either go through the entire pipeline or directly start at any stage. Gradio (abid2019gradiohasslefreesharingtesting) is an open-sourced Python package that allows developers to quickly build a web demo or application for machine learning. LlamaBoard is the user-friendly GUI (Graphical User Interface) of LlamaFactory (zheng-etal-2024-llamafactory). The GUI of our system is powered by Gradio and draws inspiration from the design of LlamaBoard.
@@ 3.1 Construct Victim Models
When a new language model B emerges, in order to better evaluate the adversarial robustness of B, we need to quantitatively and thoroughly perform evaluation on multiple downstream tasks. For the purpose of stress-testing the adversarial robustness of B more effectively, i.e., constructing a stronger adversarial robustness benchmark with high quality, we can choose at least one previous SOTA or similar-structured language model A to implement textual adversarial attacks on it to generate updated adversarial texts. We can also follow this stage when a new downstream dataset n is available.
@@ 3.2 Generate Adversarial Examples
The way human authors keep thinking and writing adversarial texts (wallace-etal-2019-trick) is high-labor-cost. With the emergence of automated textual adversarial attacks, such as TextFooler (DBLP:conf/aaai/JinJZS20), BERT-ATTACK (li-etal-2020-bert-attack), SemAttack (wang-etal-2022-semattack), and TextCheater (10345721), adversarial text generation has become relatively easy. We can directly enter this stage when a new textual adversarial attack N appears.
In this stage, we implement textual adversarial attacks I,II,...,N on the victim language models constructed from language model A upon the test set of downstream datasets 1,2,...,n to generate the first-round adversarial texts automatically. The adversarial example generation stage is depicted in the second part of Figure 2.
@@ 3.3 Construct High-Quality Benchmarks
The construction process of AdvGLUE (NEURIPS_DATASETS_AND_BENCHMARKS2021_335f5352), a widely used adversarial robustness benchmark, tells us that most textual adversarial attack methods are prone to generating invalid or ambiguous adversarial texts, with around 90% either changing the original semantics or hindering the annotators’ unanimity. Therefore, human annotation is indispensable and can make benchmarks more practical and relevant. In order to reduce the cost of human annotation, the first-round adversarial texts need to be screened automatically first using appropriate filter conditions. Due to the fact that humans perceive texts through their eyes and brains, both filter conditions and human annotation should follow the visual and semantic similarity between adversarial texts and original texts. Filter conditions can be the following metrics: Edit Distance, Normalized Cross-Correlation Coefficient (from the perspective of visual similarity); Cosine Similarity, BERTScore (DBLP:conf/iclr/ZhangKWWA20) (from the perspective of semantic similarity); and so on. Human annotation still requires additional consideration of annotators’ unanimity so that adversarial texts can be deemed human-acceptable. For example, given an original text and an adversarial text, we ask several annotators to score the human acceptance of the adversarial text based on the visual and semantic similarity between the two texts, from 1 to 5. The higher the score, the higher the human acceptance. If all annotators score the human acceptance of the adversarial text as 4 or 5, the adversarial text will be included in the adversarial robustness benchmark.
@@ 3.4 Evaluate Adversarial Robustness
This stage aims at quantitatively and thoroughly evaluating the adversarial robustness of new language models using the constructed high-quality adversarial robustness benchmark.
The adversarial robustness benchmark x is a collection of $n$ subsets, each of which contains high-quality adversarial texts generated from the test set of the corresponding downstream dataset. We take the average accuracy on $n$ subsets as the adversarial robustness ($AdvRobust$) of the new language model B on x, denoted as:
@@
In this stage, we utilize the constructed high-quality adversarial robustness benchmark x to evaluate the adversarial robustness of the language model B quantitatively and thoroughly. The adversarial robustness evaluation stage is depicted in the fourth part of Figure 2.